How many possible samples n=2 can be drawn from a population of size 14?

1
Expert's answer
2022-05-12T07:28:02-0400

The number of possible samples which can be selected without replacement is

"_NC_n=\cfrac{N! } {n! \cdot(N-n)! };"

"_{14}C_2=\cfrac{14! } {2! \cdot(14-2)! }=\cfrac{13\cdot14}{2}=91."